Use data from the following table which shows preferences for various pets by owner gender.
Gender of owner Cats Dogs Other Pets Total
Female 100 50 30 180
Male 50 50 20 120
Total 150 100 50 300
Part II: Find the probability that a randomly chosen person has the following characteristics:
(a) Is female and owns a dog, P(F  D)
(b) Is male and owns a dog, P(M  D)
[5 marks]


Part II: Find the following conditional probabilities for a randomly chosen person:
(c) Owns a dog, given that the person is female, P(D|F)
[5 marks]

(d) Owns a dog, given that the person is male, P(D|M)

Answer by stanbon(75887)

Use data from the following table which shows preferences for various pets by owner gender.
Gender of owner Cats Dogs Other Pets Total
Female 100 50 30 180
Male 50 50 20 120
Total 150 100 50 300
Part II: Find the probability that a randomly chosen person has the following characteristics:
(a) Is female and owns a dog, P(F | D) = 50/100 = 1/2
(b) Is male and owns a dog, P(M | D) = 50/100 = 1/2
-------------------------------------------------------------------
Part II: Find the following conditional probabilities for a randomly chosen person:
(c) Owns a dog, given that the person is female, P(D|F) = 50/180
(d) Owns a dog, given that the person is male, P(D|M) = 50/120
===================
Cheers,
Stan H.
=========================